# Measuring quantum relative entropy with finite-size effect

### v1

@@ Abstract
We study the estimation of relative entropy $D(\rho\|\sigma)$ when $\sigma$ is known. We show that the Cramér-Rao type bound equals the relative varentropy. Our estimator attains the Cramér-Rao type bound when the dimension $d$ is fixed. It also achieves the sample complexity $O(d^{2})$ when the dimension $d$ increases. This sample complexity is optimal when $\sigma$ is the complexity mixed state. Also, it has time complexity $O(d^{5}\mathop{\rm polylog}\nolimits d)$. Our proposed estimator unifiedly works under both settings.
@@ I Introduction
The estimation of $D(\rho\|\sigma)$ with the unknown state $\rho$ is formulated as a special case of the estimation of a function output with nuisance parameter, which was formulated in [14, Section 4.3]. In this formulation, the optimal performance of the estimation precision is given by the Cramér-Rao type bound with nuisance parameters, which equals the inverse of the projected symmetric logarithmic derivative (SLD) Fisher information in the one-parameter case. As the first step, using the results for estimation with nuisance parameters, we show that the Cramér-Rao type bound of this problem equals the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which appears in the second-order theory of hypothesis testing. The meaning of the Cramér-Rao type bound is the following; when $n$ copies of the unknown state $\rho$ is prepared, the mean square of our best estimator behaves as $\frac{V(\rho\|\sigma)}{n}+o(\frac{1}{n})$. Under special models, the recent paper [15] showed the existence of an estimator whose $n$ times the mean square error equals the Cramér-Rao type bound without nuisance parameter. However, such an estimator does not exist in general.
@@
In practice, it is necessary to find an estimator to achieve the above performance. In addition, it is also desired to clearly evaluate the difference between its mean square error and the Cramér-Rao type bound. Since our problem is one-parameter estimation, it is possible to achieve the above bound without the use of collective measurements across several quantum systems. However, this method requires an adaptive choice of measurement. One typical method for this approach is a two-step estimator, which is composed of the pre-estimation stage and the local optimal estimation. When we have $n$ copies of the unknown state $\rho$, theoretically, in order to achieve the optimal bound, it is sufficient to allocate $\sqrt{n}$ copies to the pre-estimation stage [17, 16]. Although this method achieves the optimal mean square error in the order $\frac{1}{n}$, i.e., the first order, it is not so clear how the mean square error is close to the leading term $\frac{V(\rho\|\sigma)}{n}$. From the practical viewpoint, it is needed to upper bound this difference by using the number $n$ of unknown copies. Unfortunately, the above method has difficulty to evaluate the difference between its mean square error and $\frac{V(\rho\|\sigma)}{n}$. The papers [18, 19] also proposed different adaptive methods for the parameter estimation, but their methods have the same problem. In addition, the above adaptive approach does not work when the dimension $d$ of the quantum system increases in addition to the number $n$ of copies.
@@
As the second step, to satisfy the above requirements, we propose an alternative estimation method by using Schur transform [20]. Then, we evaluate its mean square error and clarify its difference from the Cramér-Rao type bound with a finite number $n$. In addition, we evaluate the tail probability.
As the third step, similar to the papers [6, 7], we study the case when the dimension $d$ increases when $n$ increases. Unlike the estimation of von Neumann entropy, the range of the quantum relative entropy nor the relative varentropy is not bounded even with fixed dimension $d$. Therefore, to study the sample complexity, we need to restrict the possible range of $\rho$ and $\sigma$. We show that the sample complexity of our method is $O(d^{2})$ under a certain natural restriction for this range. In particular, the restriction of the possible range of $\sigma$ is crucial in this analysis. We also show that our time complexity is $O(d^{5}\mathop{\rm polylog}\nolimits d)$.
When $\sigma$ is fixed to the completely mixed state, our problem is the same as the estimation of von Neumann entropy, which was studied in the papers [6, 7]. Our result contains the case studied in the papers [6, 7]. In this special case, the paper [6] also showed that the lower bound of the sample complexity is $O(d^{2})$ when our measurement is limited to Schur sampling. The paper [7] also showed that the lower bound of the sample complexity is $O(d)$ without this restriction. In this special case, we show that our measurement can be restricted to Schur sampling so that the lower bound by [6] holds even without this restriction.
The remaining part of this paper is organized as follows. First, Section II derives the Cramér-Rao type bound for the estimation of quantum relative entropy. Section III presents our estimator for quantum relative entropy, and its performance. Then, this section shows that this estimator asymptotically attains the Cramér-Rao type bound when $d$ is fixed. Section IV studies the sample complexity of our method when $d$ increases. Section V studies the time complexity of our method when $d$ increases. Section VI shows the optimality of the use of Schur sampling when $\sigma$ is fixed to the completely mixed state. Appendices are devoted to the proofs of the results presented in Section III.
@@ II Cramér-Rao type bound
To clarify the minimum mean square error with the first order $\frac{1}{n}$, we calculate the Cramér-Rao type bound with nuisance parameters of $D(\rho\|\sigma)$ at $\rho=\rho_{*}$. For this aim, we review the Cramér-Rao type bound for a general function $f(\rho)$ in the full parameter model under a $d$-dimensional system as follows. We choose a parametrization $\rho_{\theta}$ with $\theta=(\theta^{1},\theta^{2},\ldots,\theta^{d^{2}-1})\in\mathbb{R}^{d^{2}-1}$ such that the first parameter gives the value of the function to be estimated and the other parameters are orthogonal to the first parameter at $\rho=\rho_{*}$. That is, our parameter needs to satisfy the following conditions.
@@
where $(a)$ follows from [21, Eq. (6.32)].
@@
which implies the condition (5). Therefore, since all the conditions (1) – (5) hold, due to Proposition 1, the Cramér-Rao type bound for the estimation of $D(\rho\|\sigma)$ is $\langle L_{1},L_{1}\rangle_{\rho_{*}}^{-1}=V(\rho_{*}\|\sigma)^{-1}$ at $\theta=\theta_{0}$. That is, the minimum mean square error under the local unbiasedness condition is $V(\rho_{*}\|\sigma)$. That is, the Cramér-Rao type bound for the estimation of $D(\rho\|\sigma)$ is the relative varentropy $V(\rho_{*}\|\sigma)$.
@@ III Estimator based on Schur transform
Similar to the papers [22, 3, 8, 9, 10, 11, 12, 13, 4, 23, 24], we employ Schur duality of ${\cal H}^{\otimes n}$. For this aim, we prepare several notations. A sequence of monotone-increasing non-negative integers $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{d})$ is called Young index. Although many references [24, 26, 25] define Young index as monotone-decreasing non-negative integers, we define it in the opposite way for notational convenience. We denote the set of Young indices $\bm{\lambda}$ with condition $\sum_{j=1}^{d}\lambda_{j}=n$ by $Y_{d}^{n}$ Then, as explained in [24, Section 6.2], we have
@@
To evaluate the mean square error of this estimator, we denote the dimension of $\oplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}$ by $d_{n}$. Since as shown in [24, (6.16) and (6.18)], we have
@@
$\leq$ $\Big{(}\frac{1}{\sqrt{n}}\sqrt{V(\rho\|\sigma)}+\frac{1}{n}\log d_{n}\Big{)}^{2}.$ (21)
This theorem is proved in Appendix A. The difference between the upper bound (21) and the leading term $\frac{1}{n}V(\rho\|\sigma)$ is $\frac{1}{n^{2}}\log d_{n}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\log d_{n}$. Using (20), this different is upper bounded by $\frac{1}{n^{2}}(\frac{(d+2)(d-1)}{2}\log(n+1))^{2}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\frac{(d+2)(d-1)}{2}\log(n+1)$. $n$ times of these terms go to zero.
@@
$\delta_{\rho,\sigma}^{(n),+,R-D(\rho\|\sigma)}\leq$ $d_{n}^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)},$ (26)
$\delta_{\rho,\sigma}^{(n),-,R-D(\rho\|\sigma)}\leq$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}+d_{n}e^{-nr}.$ (27)
@@ IV Analysis with increasing $d$
Similar to [6, 7], we consider the case when the dimension $d$ increases. When $n=O(d^{4+\delta})$, $\frac{1}{\sqrt{n}}\log d_{n}\leq\frac{1}{\sqrt{n}}\frac{(d+1)(d-1)}{2}\log(n+1)\to 0$. In this case, due to Theorem 2, the Cramér-Rao type bound can be attained because the second term of RHS in (21) goes to zero.
@@
For example, the condition (28) holds under certain conditions.
@@
When we choose $c$ to be $\frac{(\sqrt{c_{0}}+4)^{2}}{\epsilon^{\prime}\epsilon^{2}}$ for any $\epsilon,\epsilon^{\prime}>0$, we have
@@
Hence, we obtain an upper bound $O(d^{2})$ of the sample complexity for the estimation of $D(\rho\|\sigma_{d})$ when $\rho$ belongs to ${\cal S}_{d}$.
Proof.
We choose $s\in(0,1)$. Using Weyl’s dimension formula [25, Theorem 7.1.7], we have
@@
Since the term $\frac{1}{n}(d-1)\log(n+1)$ goes to zero, the term $\frac{1}{n}\log d_{n}$ is evaluated as
$\lim_{n\to\infty}\frac{1}{n}\log d_{n}\leq\min_{0<s<1}\frac{c^{s-1}}{s(1-s)}.$ (38)
@@
$\lim_{n\to\infty}\frac{1}{\sqrt{n}}\sqrt{V(\rho\|\sigma)}=\frac{\sqrt{c_{0}}}{\sqrt{c}}.$ (39)
@@
$\lim_{n\to\infty}\mathop{\rm MSE}\nolimits_{n}(\rho\|\sigma_{d})$
@@
The key point of this proof is the derivation (34). In particular, the derivation (35) is the heart in this part. The idea of this derivation comes from the proof of (12) given in Appendix IV of the paper [27], which gives the security analysis in information theoretic security.

### v2

@@ Abstract
We study the estimation of relative entropy $D(\rho\|\sigma)$ when $\sigma$ is known. We show that the Cramér-Rao type bound equals the relative varentropy. Our estimator attains the Cramér-Rao type bound when the dimension $d$ is fixed. It also achieves the sample complexity $O(d^{2})$ when the dimension $d$ increases. This sample complexity is optimal when $\sigma$ is the completely mixed state. Also, it has time complexity $O(d^{6}\mathop{\rm polylog}\nolimits d)$. Our proposed estimator unifiedly works under both settings.
relative entropy; sample complexity; Cramer-Rao type bound; Schur composition
@@ I Introduction
The estimation of $D(\rho\|\sigma)$ with the unknown state $\rho$ is formulated as a special case of the estimation of a function output with nuisance parameters. This general problem setting has been studied in various researches [14, 15, 16, 17, 18, 19]. The reference [19, Section III] addressed the estimation problem of $D(\rho\|\sigma)$ as a special case, and showed that the estimation error is lower bounded by the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which appears in the second-order theory of hypothesis testing [20, 21] while it did not show its attainability. Under a general setting, the reference [16, Section 4.3] showed that the optimal performance of the estimation precision is given by the Cramér-Rao type bound $C$ with nuisance parameters, which equals the inverse of the projected symmetric logarithmic derivative (SLD) Fisher information in the one-parameter case. The meaning of the Cramér-Rao type bound $C$ is the following; when $n$ copies of the unknown state $\rho$ is prepared under a fixed model, the mean square of our best estimator behaves as $\frac{C}{n}+o(\frac{1}{n})$. When we have no nuisance parameter, the recent paper [22] showed the existence of an estimator whose mean square error (MSE) equals $\frac{C}{n}$ without additional term under special models. However, such an estimator does not exist in general.
@@
In this paper, as the first step, using the results for estimation with nuisance parameters in [16, Section 4.3], we show that the Cramér-Rao type bound $C$ of this problem equals the relative varentropy $V(\rho\|\sigma):=\mathrm{Tr}\rho((\log\rho-\log\sigma)-D(\rho\|\sigma))^{2}$, which was obtained as a lower bound in [19]. In practice, it is necessary to find an estimator to achieve the MSE $\frac{C}{n}+o(\frac{1}{n})$. In addition, it is also desired to clearly evaluate the difference between its mean square error and the Cramér-Rao type bound $\frac{C}{n}$. Since our problem is one-parameter estimation, it is possible to achieve the above bound without the use of collective measurements across several quantum systems. However, this method requires an adaptive choice of measurement. One typical method for this approach is a two-step estimator, which is composed of the pre-estimation stage and the local optimal estimation. When we have $n$ copies of the unknown state $\rho$, theoretically, in order to achieve the optimal bound, it is sufficient to allocate $\sqrt{n}$ copies to the pre-estimation stage [23, 15]. Although this method achieves the optimal mean square error in the order $\frac{1}{n}$, i.e., the first order, it is not so clear how the mean square error is close to the leading term $\frac{C}{n}$. From the practical viewpoint, it is necessary to upper bound this difference by using the number $n$ of unknown copies. Unfortunately, the above method has difficulty to evaluate the difference between its mean square error and $\frac{C}{n}$. The papers [24, 25] also proposed different adaptive methods for parameter estimation, but their methods have the same problem. In addition, the above adaptive approach does not work when the dimension $d$ of the quantum system increases in addition to the number $n$ of copies.
@@
As the second step of this paper, to satisfy the above requirements, we propose an alternative estimation method by using Schur transform [26]. Then, we evaluate its mean square error and clarify its difference from the Cramér-Rao type bound with a finite number $n$. In addition, we evaluate the tail probability.
As the third step, similar to the papers [6, 7], we study the case when the dimension $d$ increases when $n$ increases. Unlike the estimation of von Neumann entropy, the range of the quantum relative entropy nor the relative varentropy is not bounded even with fixed dimension $d$. Therefore, to study the sample complexity, we need to restrict the possible range of $\rho$ and $\sigma$. We show that the sample complexity of our method is $O(d^{2})$ under a certain natural restriction for this range. Our method extensively improves the sample complexity over the simple application of the full tomography [27]. In particular, the restriction of the possible range of $\sigma$ is crucial in this analysis. We also show that our time complexity is $O(d^{6}\mathop{\rm polylog}\nolimits d)$.
When $\sigma$ is fixed to the completely mixed state, our problem is the same as the estimation of von Neumann entropy, which was studied in the papers [6, 7]. Our result contains the case studied in the papers [6, 7]. In this special case, the paper [6] also showed that the lower bound of the sample complexity is $O(d^{2})$ for the Empirical Young Diagram (EYD) algorithm, which is given as the entropy of the diagonal state determined by the Young indecies. Although this case improves the sample complexity over the simple application of the full tomography, the improvement is incremental. The paper [7] also showed that the lower bound of the sample complexity is $O(d)$ without this restriction. In this special case, we show that our measurement can be restricted to Schur sampling.
The remaining part of this paper is organized as follows. First, Section II derives the Cramér-Rao type bound for the estimation of quantum relative entropy. Section III presents our estimator for quantum relative entropy, and its performance. Then, this section shows that this estimator asymptotically attains the Cramér-Rao type bound when $d$ is fixed. Section IV studies the sample complexity of our method when $d$ increases. Section V studies the time complexity of our method when $d$ increases. Appendices are devoted to the proofs of the results presented in Section III.
@@ II Cramér-Rao type bound
To clarify the minimum mean square error with the first order $\frac{1}{n}$, we discuss the Cramér-Rao type bound $C$ with nuisance parameters, which is mathematically formulated as a lower bound of MES under the local unbiasedness condition. The reference [19, (3.24), (3.52)] showed that the varentropy $V(\rho_{*}\|\sigma)$ is a lower bound under the estimation of $D(\rho\|\sigma)$ at $\rho=\rho_{*}$. That is, they showed that
$C\geq V(\rho_{*}\|\sigma),$ (1)
and did not discuss whether the equality holds.
In the following, we show the equality in (1). For this aim, we review the Cramér-Rao type bound $C$ for a general function $f(\rho)$ in the full parameter model under a $d$-dimensional system as follows. We choose a parametrization $\rho_{\theta}$ with $\theta=(\theta^{1},\theta^{2},\ldots,\theta^{d^{2}-1})\in\mathbb{R}^{d^{2}-1}$ such that the first parameter gives the value of the function to be estimated and the other parameters are orthogonal to the first parameter at $\rho=\rho_{*}$. That is, our parameter needs to satisfy the following conditions.
@@
where $(a)$ follows from [28, Eq. (6.32)].
@@
which implies the condition (6). Therefore, since all the conditions (2) – (6) hold, due to Proposition 1, the Cramér-Rao type bound for the estimation of $D(\rho\|\sigma)$ is $\langle L_{1},L_{1}\rangle_{\rho_{*}}^{-1}=V(\rho_{*}\|\sigma)^{-1}$ at $\theta=\theta_{0}$. That is, the minimum mean square error under the local unbiasedness condition is $V(\rho_{*}\|\sigma)$. That is, the Cramér-Rao type bound for the estimation of $D(\rho\|\sigma)$ is the relative varentropy $V(\rho_{*}\|\sigma)$. Thus, the quality in (1) holds.
@@ III Estimator based on Schur transform
Similar to the papers [29, 3, 8, 9, 10, 11, 12, 13, 4, 30, 31], we employ Schur duality of ${\cal H}^{\otimes n}$. For this aim, we prepare several notations. A sequence of monotone-increasing non-negative integers $\bm{\lambda}=(\lambda_{1},\ldots,\lambda_{d})$ is called Young index. Although many references [31, 33, 32] define Young index as monotone-decreasing non-negative integers, we define it in the opposite way for notational convenience. We denote the set of Young indices $\bm{\lambda}$ with the condition $\sum_{j=1}^{d}\lambda_{j}=n$ by $Y_{d}^{n}$ Then, as explained in [31, Section 6.2], we have
@@
To evaluate the mean square error of this estimator, we denote the dimension of $\oplus_{\bm{\lambda}\in Y_{d}^{n}}{\cal U}_{\bm{\lambda}}$ by $d_{n,d}$. Since as shown in [31, (6.16) and (6.18)], we have
@@
$\leq$ $\Big{(}\frac{1}{\sqrt{n}}\sqrt{V(\rho\|\sigma)}+\frac{1}{n}\log d_{n,d}\Big{)}^{2}.$ (22)
This theorem is proved in Appendix A. The difference between the upper bound (22) and the leading term $\frac{1}{n}V(\rho\|\sigma)$ is $\frac{1}{n^{2}}\log d_{n,d}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\log d_{n,d}$. Using (21), this different is upper bounded by $\frac{1}{n^{2}}(\frac{(d+2)(d-1)}{2}\log(n+1))^{2}+\frac{2}{n\sqrt{n}}\sqrt{V(\rho\|\sigma)}\frac{(d+2)(d-1)}{2}\log(n+1)$. $n$ times of these terms go to zero.
@@
$\delta_{\rho,\sigma}^{(n),+,R-D(\rho\|\sigma)}\leq$ $d_{n,d}^{\alpha}e^{-n\alpha(D_{1-\alpha}(\rho\|\sigma)-R)},$ (27)
$\delta_{\rho,\sigma}^{(n),-,R-D(\rho\|\sigma)}\leq$ $e^{-n(R-r-\alpha(D_{1+\alpha}(\rho\|\sigma))}+d_{n,d}e^{-nr}.$ (28)
@@ IV Analysis with increasing $d$
Similar to [6, 7], we consider the case when the dimension $d$ increases. When $n=O(d^{4+\delta})$, $\frac{1}{\sqrt{n}}\log d_{n,d}\leq\frac{1}{\sqrt{n}}\frac{(d+1)(d-1)}{2}\log(n+1)\to 0$. In this case, due to Theorem 2, the Cramér-Rao type bound can be attained because the second term of RHS in (22) goes to zero.
@@
For example, when $\sigma_{d}$ is fixed to the completely mixed state $\rho_{\mathop{\rm mix}\nolimits}$, $V(\rho\|\rho_{\mathop{\rm mix}\nolimits})$ equals the varentropy $\mathrm{Tr}\rho(\log\rho)^{2}-(\mathrm{Tr}\rho\log\rho)^{2}$. Since Lemma 8 of [3] showed that $\mathrm{Tr}\rho(\log\rho)^{2}\leq(\log d)^{2}$, the condition (29) holds even with ${\cal S}_{d}={\cal S}({\cal H}_{d})$ so that the term $\frac{1}{n}\log d_{n,d}$ is the major term in the RHS of (22). Further, the condition (29) holds with ${\cal S}_{d}={\cal S}({\cal H}_{d})$ under certain conditions for $\sigma_{d}$.
@@
When we choose $c$ to be $\frac{(\sqrt{c_{0}}+4)^{2}}{\epsilon^{\prime}\epsilon^{2}}$ for any $\epsilon,\epsilon^{\prime}>0$, i.e., $n=\frac{(\sqrt{c_{0}}+4)^{2}}{\epsilon^{\prime}\epsilon^{2}}d^{2}$, we have
@@
Hence, when the condition (34) is imposed for the estimation error, we obtain an upper bound $O(d^{2})$ of the sample complexity for the estimation of $D(\rho\|\sigma_{d})$ when $\rho$ belongs to ${\cal S}_{d}$.
Now, we compare the sample complexity of our method with the sample complexity of the simple application of full tomography [27, 34]. In particular, we employ the error analysis on the trance norm [27] rather than other error criteria because the trance norm is most useful to evaluate the error of the relative entropy. Their method achieves the estimation error
$\|\hat{\rho}-\rho\|_{1}\leq\epsilon$ (35)
with high probability by sample complexity $O(d^{2}/\epsilon^{2})$, where $\hat{\rho}$ is the estimated density and $\rho$ is the true density. When the known density $\sigma_{d}$ is the completely mixed state, our problem is reduced to the estimation of von Neumann entropy $S(\rho)$. Fannes inequality implies the relation
$|S(\hat{\rho})-S(\rho)|\leq\|\hat{\rho}-\rho\|_{1}\log d.$ (36)
That is, to achieve the condition
$|S(\hat{\rho})-S(\rho)|\leq\epsilon^{\prime},$ (37)
we need the sample complexity $O(d^{2}(\log d)^{2}/{\epsilon^{\prime}}^{2})$. Although our method and the method by [6] improves the sample complexity over the above method, the degree of the improvement is incremental in this case.
When the known state $\sigma_{d}$ is a general state to satisfy the condition of Lemma 5, i.e., the minimum eigenvalue of $\sigma_{d}$ is lower bounded by $e^{-td}$ with a fixed real number $t>0$, the part $\log\sigma_{d}$ has the operator norm $td$ at most. Hence, we have
$|D(\hat{\rho}\|\sigma_{d})-D(\rho\|\sigma_{d})|$
$\leq$ $|S(\hat{\rho})-S(\rho)|+|\mathrm{Tr}(\hat{\rho}-\rho)\log\sigma_{d}|$
$\leq$ $\|\hat{\rho}-\rho\|_{1}\log d+\|\hat{\rho}-\rho\|_{1}\|\log\sigma_{d}\|$
$\leq$ $\|\hat{\rho}-\rho\|_{1}(\log d+td).$ (38)
That is, to achieve the condition
$|D(\hat{\rho}\|\sigma_{d})-D(\rho\|\sigma_{d})|\leq\epsilon^{\prime},$ (39)
we need the sample complexity $O(d^{2}(\log d+td)^{2}/{\epsilon^{\prime}}^{2})=O(d^{4}/{\epsilon^{\prime}}^{2})$. Our method greatly improves the sample complexity over the the simple application of full tomography [27]. That is, our extension to the estimation of the relative entropy realizes significant improvement in the sample complexity over the above method.
Proof of Theorem 6.
We choose $s\in(0,1)$. Using Weyl’s dimension formula [32, Theorem 7.1.7], we have
@@
Since the term $\frac{1}{n}(d-1)\log(n+1)$ goes to zero, the term $\frac{1}{n}\log d_{n,d}$ is evaluated as
$\lim_{d\to\infty}\frac{1}{n}\log d_{n,d}\leq\min_{0<s<1}\frac{c^{s-1}}{s(1-s)}.$ (44)
@@
$\lim_{d\to\infty}\frac{1}{\sqrt{n}}\max_{\rho\in{\cal S}_{d}}\sqrt{V(\rho\|\sigma_{d})}=\frac{\sqrt{c_{0}}}{\sqrt{c}}.$ (45)
@@
$\lim_{d\to\infty}\max_{\rho\in{\cal S}_{d}}\mathop{\rm MSE}\nolimits_{n}(\rho\|\sigma_{d})$
@@
The key point of this proof is the derivation (40). In particular, the derivation (41) is the heart in this part. The idea of this derivation comes from the proof of (12) given in Appendix IV of the paper [35], which gives the security analysis in information-theoretic security.